Alvin paddled for 5 hours with a 2-km/h current to reach a campsite. The return trip against the same current took 7 hours. What is the speed of the boat in still water?
Answer by mananth(16946)

Boat speed =x kmph
current speed 2 kmph

Time with current= 5 hours
time against current 7 hours
Distance = same d=rt
5(x+2)= 7(x-2)
5x+10=7 x-14
7x-5x = 10+ 14
2 x= 24
x = 12 kmph Boat speed